Inspect the following without finding the determinant:

1.

[1 0 0 0]

[0 1 0 0]

[0 0 1 0]

[0 0 0 - 2]


2. [1 0 0 0]

[0 1 0 0]

[0 0 0 1]

[0 0 1/4 0]


1
Expert's answer
2021-06-07T14:37:50-0400

If all elements of a row (or column) of a determinant are multiplied by some scalar number k, the value of the new determinant is k times of the given determinant.


If any two row (or two column) of a determinant are interchanged the value of the determinant is multiplied by "-1."


In identity matrix "a_{ii}=1," and all other elements = 0, hence the determinant is 1.


1.


"\begin{vmatrix}
 1 & 0 & 0 & 0 \\
 0 & 1 & 0 & 0 \\
 0 & 0 & 1 & 0 \\
 0 & 0 & 0 & -2 \\
\end{vmatrix}=-2\begin{vmatrix}
 1 & 0 & 0 & 0 \\
 0 & 1 & 0 & 0 \\
 0 & 0 & 1 & 0 \\
 0 & 0 & 0 & 1 \\
\end{vmatrix}=-2(1)=-2"

2.


"\begin{vmatrix}
 1 & 0 & 0 & 0 \\
 0 & 1 & 0 & 0 \\
 0 & 0 & 0 & 1 \\
 0 & 0 & 1/4 & 1 \\
\end{vmatrix}=1/4\begin{vmatrix}
 1 & 0 & 0 & 0 \\
 0 & 1 & 0 & 0 \\
 0 & 0 & 0 & 1 \\
 0 & 0 & 1 & 1 \\
\end{vmatrix}"


"=-1/4\begin{vmatrix}
 1 & 0 & 0 & 0 \\
 0 & 1 & 0 & 0 \\
 0 & 0 & 1 & 0 \\
 0 & 0 & 0 & 1 \\
\end{vmatrix}=-1/4"